1- A number of publicly traded firms pay no dividends yet investors are willing to buy
shares in these firms. How is this possible? Does this violate our basic principle of stock
valuation? Explain your answer.


2- The risk free rate of return is 8 percent; the expected rate of return on the market is 12
percent. Stock X has a beta coefficient of 1.3, an earnings and dividend growth rate of 7
percent, and a current dividend of RM2.40. If the stock is selling for RM35, what should
you do?
1
Expert's answer
2016-05-12T08:36:02-0400
1- If a number of publicly traded firms pay no dividends yet investors are willing to buy
shares in these firms, then this violates our basic principle of stock valuation, because to calculate the current market price of stock we need to know the amount of dividends paid.

2- Rf = 8%; r = 12%, b = 1.3, g = 7%, Do = RM2.40. In this case the market price of the stock should be P = Do*(1 + g)/(r - g) = 2.4*1.07/(0.12 - 0.07) = RM51.36.
If the stock is selling for RM35, then you shouldn't sell (or should buy) it, because its price is undervalued.